Jack is 7 years older than Conor
(a) If jacks age is x years what is Conor’s age?
(b) If twice the sun of their ages is 66 find Conor’s age

Answer by ikleyn(52797)

.

(a) x-7 years.      ANSWER


(b)  2*(x +  (x-7)) = 66


     x + (x-7)      = 33


     2x - 7         = 33


     2x             = 33 + 7 = 40


      x                      = 40/2 = 20.   (Jack's age).


ANSWER.  Connor is 20-7 = 13 years old.

Solved.